A heater is being designed that uses a coil of 14-gauge (1.628 mm diameter) manganin wire to generate 301 W using a voltage of V = 118 V. How long should the engineer make the
Answer #1

Given d = 1.628 mm, P = 301 W, V = 118 V.

Radius r = d/2 = 1.628/2 = 0.814 mm